If the reserve ratio is 100-percent, then a new deposit of $500 into a bank account

eventually increases the money supply by $500

leaves the size of the money supply unchanged

eventually decreases the size of the money supply by $500.

None of the above is correct.
Answer #1

The correct option is B leaves the size of the money supply unchanged

Mathematically, the monetary multiplier would be 1/RR = 1/1 = 1, so the amount of new money would be $500 × MM – $500 = $500 × 1 – $500 = $0